1) Sunset rents an SUV at $21.95 plus $0.23 per mile. Sunrise rents the same vehicle for $24.95 plus $0.19 per mile. For what mileage is the cost the same?
2) Your piggy bank has 25 coins in it; some are quarters and some are nickels. You have $3.45. How many nickels do you have?
3) There were 1500 people at a high school football game. Student tickets were $2.00 and adult tickets were $3.50. The total receipts for the game were $3825. How many students bought tickets?

Answer by stanbon(75887)

1. Let number of miles driven be "x", then:
Sunset cost is 2195 + 23x cents for x miles driven
Sunrise cost is 2495 +19x cents for x miles driven
EQUATION: Set costs equal to one another and solve
for "x".
2. Let number of nickels be "x".
Then number of quarter is "25-x".
Value of nickels is (5)x cents
Value of quarters is (25)(25-x)= 625-25x cents
EQUATION:
Value of nickels + Value of Quarters is 345 cents
5x + 625-25x = 345
-20x = -280
x = 14 (number of nickels)
25-x = 11 (number of quarters)
#3 This problem is done the same way as #2 except now
the coins are tickets and their respective values are
200 and 350 cents. Give it a try.
Cheers,
Stan H.